# Elastic tensor-derived properties of composition-dependent disordered refractory binary alloys

### v2

@@ Abstract
The elastic tensor provides valuable insight into the mechanical behavior of a material with lattice strain, such as disordered binary alloys. Density functional perturbation theory (DFPT) based on density functional theory (DFT) provides a powerful mechanism for computing and probing the microscopic features of elastic tensor-related properties. Here we present results for the rigid-ion and relaxed-ion elastic tensors computed using DFPT, for a comprehensive set of structural refractory body-centered cubic (BCC) binary alloys of molybdenum (Mo), niobium (Nb), tantalum (Ta), and tungsten (W). For the first time, we have mapped the heterogeneity in elastic constants and shear modulus and associated relaxation fields at each lattice site by computing the force response internal strain tensor ($\Lambda$) and displacement response internal strain tensors ($\Gamma$). Derived properties—the bulk modulus ($B$), shear modulus ($G$), Young’s modulus ($E$), Poisson’s ratio ($\nu$), Pugh’s ratio ($B/G$), Cauchy pressure and elastic anisotropy—are reported as a function of composition for all refractory binaries. The computed mechanical properties data for the refractory binary alloys at systematically-varied Mo, Nb, Ta, and W compositions are in excellent agreement with available experimental data.
@@
Refractory alloy , Density functional theory , Compositionally-complex alloys , Elastic properties , Property anisotropy
journal: Comp. Mat. Sci.
[…]
[1]organization=Department of Mechanical Engineering, addressline=University of New Mexico, postcode=87131, city=Albuquerque, country=USA
[…]
[2]organization=Department of Chemistry and Chemical Biology, Department of Physics and Astronomy, and Center for Quantum Information and Control, addressline=University of New Mexico, postcode=87131, city=Albuquerque, country=USA
[…]
[3]organization=Department of Mechanical Engineering, addressline=University of New Mexico, postcode=87131, city=Albuquerque, country=USA
@@ 1 Introduction
The elastic tensor [1] and its elastic constant components give a complete description, in terms of deformation, of material response under external stress within the elastic limit. Inducing lattice strain through alloying is a key design strategy for tuning and optimizing the properties of compositionally-complex materials [2, 3] such as multi-component refractory alloys comprised of four or more principal elements and alloying elements [4]. The elastic tensor can therefore serve as a valuable quantitative probe of elastic strain effects associated with changes in chemical composition, and their impact on mechanical [5, 6, 7, 8, 9], electronic [10, 11], thermodynamic [12, 13], and structural [14, 15, 16] properties in these materials.
@@
Refractory alloys show great potential to meet the demands for high-temperature materials due to their high melting points and ability to maintain mechanical properties at temperatures higher than current Ni-based superalloys [17]. The refractory elements Nb, Ta, Mo, and W of groups V and VI comprise the BCC structural refractory metals family, characterized by high mechanical strength and melting temperatures [18]. Their structural refractory binary alloys MoNb, MoTa, MoW, NbTa, WNb and WTa form solid solutions of BCC crystal structure across their entire compositional space [19]. Elastic constant measurements are typically reported only for a limited number of alloy classes and compositions, and experiments often exhibit substantial disagreement even between values measured using the same technique and on the same material [20, 21]. Multiple factors can influence differences in reported experimental values, such as impurities and measurement temperature [22, 23, 24, 25]. Previous efforts have focused on the computational determination of elastic constants using density functional theory (DFT) for selected refractory binary alloys [26, 27, 28, 29, 30]. While these investigations have provided valuable insights, they were limited to a few binary systems or computed for narrow compositional range. The computational methodologies employed in these studies were primarily validated using experimental data for pure elements. However, the electronic interactions in alloy systems are inherently more complex than in their elemental counterparts. Consequently, the direct application of methods validated for pure elements to alloy systems may introduce significant uncertainties, potentially compromising the accuracy and practical relevance of the predicted elastic properties for these alloys. Here, we establish a first-principles modeling framework based on DFT [31, 32] for performing computationally-efficient studies of elastic tensor and derived properties of disordered alloys, validated for systematically-varied Mo, Nb, Ta, and W binary alloy compositions. DFT, as implemented in state-of-the-art plane wave pseudopotential periodic electronic structure codes such as Abinit [33, 34], provides direct access to the computation of energies and stresses for obtaining the elastic tensors of the materials, with only atomic structure and nuclear charges as inputs.
@@
The principal quantity to be considered is the elastic tensor and its elastic constant components. In the traditional approach, the elastic constants $C_{ij}$ are obtained by computing either the total energy or the stress of a crystal at different strains [22, 35, 36, 37, 38]. The stress-strain relation is given by:
@@
where $\bm{\sigma}$ is the $6\times 1$ stress tensor, the $C_{ij}$ are the elastic constants, and $\bm{s}$ is the $6\times 1$ strain tensor in Voigt notation [39]. To obtain all 21 independent elastic constants, stress tensors must be computed from fully-relaxed DFT calculations for each applied strain $\bm{s}_{ij}$ (total of 6 independent strain components) for 4 different strain magnitudes [40]. Including the initial geometry optimization of the structure, a total of 24+1 independent calculations are required to compute the elastic tensor [22, 41, 42]. For each structure with its applied strain, the atomic positions must be re-relaxed prior to the computation of the stresses. These methods are therefore computationally expensive, especially for low-symmetry structures. Such low-symmetry systems are found in solid solution alloys where atomic mismatch and random configurations break the symmetry of the parent structures, as demonstrated in BCC refractory binary alloys [19]. For this reason, BCC disordered alloys are often approximated as high-symmetry structures with only three independent elastic constants ($C_{11}$, $C_{12}$ and $C_{44}$), in order to reduce the number of computations required to compute the elastic tensor [5, 26, 43, 44, 45]. Such approximations may introduce significant errors in the prediction of elastic properties, particularly in low-symmetry disordered binary alloys where the symmetry is broken due to lattice strain introduced by alloying elements.
Use of such simplifying approximations can result in the loss of critical information, such as the directional dependence of elastic properties (elastic anisotropy), which requires the computation of the complete elastic tensor. Anisotropy not only reflects the underlying atomic arrangement but also determines the material’s deformation behavior under applied stress in different crystallographic directions. Despite its significance, elastic anisotropy remains insufficiently explored across a large range of refractory binary alloy systems. The use of computational methodologies capable of accurately determining the full elastic tensor offers a pathway for systematically investigating composition-dependent elastic anisotropy in complex systems.
A faster and simpler alternative, appropriate for scaling to complex alloys, is by analytical differentiation of the energy using linear response theory [46, 47, 48, 49]. In this method, the elastic tensor is obtained by expanding the total energy with respect to infinitesimal perturbation of the system using density functional perturbation theory (DFPT) [50]. DFPT enables the computation of elastic constants without need for empirical parameters. This first-principles approach allows DFPT to scale efficiently for complex alloy systems, including those with large unit cells or reduced crystallographic symmetry, making it particularly well-suited for studying disordered or low-symmetry alloys. DFPT was initially developed for norm-conserving pseudopotentials (NCPPs) [51] and later extended to the projector augmented-wave (PAW) method by Martin et al. [52]. DFPT implemented with NCPPs has been shown to be successful in the computation of the elastic tensor for a variety of materials [53, 54, 55]. In this method, the need for a preselected list of strains as a function of material and composition is not required, which avoids the potential generation of unphysical structures for unexplored alloys. Despite its computational efficiency in determining the elastic tensor, DFPT has been applied to only a limited number of complex systems to date [56, 57, 58, 59, 60, 61, 62, 63, 64]. This work represents the first application of DFPT for computing the elastic constants of refractory binary alloys of Mo, Nb, Ta and W.
@@
Within the DFPT framework, the elastic tensor can be evaluated as the rigid-ion elastic tensor, $C_{ij}^{rigid}$, or as the relaxed-ion elastic tensor $C_{ij}^{relaxed}$ [46, 47, 48, 49]. $C_{ij}^{rigid}$ is computed with the relaxed initial structure, and strain as the perturbation; $C_{ij}^{relaxed}$ is a correction to the rigid-ion elastic tensor, with both strain and atomic displacements (nuclear relaxations) as the perturbations. In contrast to the traditional stress-strain method, each of these elastic tensor calculations requires only a single DFPT calculation, and both $C_{ij}^{rigid}$ and $C_{ij}^{relaxed}$ provide essential information about the elastic response of a material. The relaxed-ion elastic tensor is the quantity that will correspond most closely to experiment.
@@
Comparison of the rigid-ion and relaxed-ion results allows an assessment of the accuracy of the rigid-ion approximation, as well as enabling the mapping of the relaxation field of a disordered alloy. These maps can provide quantitative measures of the microscopic elastic response of the binary alloys, for later input to models of more complex multi-principal-element alloys. We define the nuclear relaxation elastic tensor $C_{ij}^{\text{nuclear}}$ as the difference between the corresponding rigid-ion and relaxed-ion quantities [49]:
@@
In the present work, we utilize DFT+DFPT to compute the single-crystal, relaxed-structure elastic tensors of the six BCC refractory binary alloys of (Mo, Nb, Ta, W)—MoNb, MoTa, MoW, NbTa, MoTa, and MoW—as a function of composition. The disordered alloys are modeled as supercells constructed using the special quasi-random structures (SQS) method [66]. The relaxed disordered structures are then used to compute the rigid-ion and relaxed-ion elastic tensors based on DFPT. We assess the effectiveness of DFPT in providing accurate elastic tensors through detailed comparison of elastic constants with available experimental data. The composition-dependent calculations serve as an important test of the ability of the SQS method to accurately model disorder in these alloys. Mechanical properties $B$, $G$, and $E$, the elastic anisotropy and $B/G$ and $C^{\prime}$ derived from the computed elastic moduli are reported as a function of composition for all refractory binaries. Computed relaxation fields for the Voigt-averaged elastic constants and shear modulus are presented for all binaries and discussed in terms of constituent element chemistry. Also, we discuss the computational efficiency of DFPT compared to traditional stress-strain methods. The in-depth understanding of trends in elastic properties and anisotropy in the disordered refractory binary alloys gained in this work through the decomposition of elastic constants into electronic and nuclear contributions is expected to contribute to new microscopic alloy computational strategies for designing refractory alloys for high-temperature applications.
@@ 2 Computational Methods
We used the open-source, periodic plane wave pseudopotential electronic structure code Abinit [33, 34] for all DFT calculations, employing the Perdew, Burke, and Ernzerhof (PBE) generalized gradient approximation (GGA) exchange-correlation functional [67], with corresponding PBE norm-conserving pseudopotentials [68]. We note that while meta-GGA methods such as r2SCAN [69] and parameterized DFT+$U$ [70] approaches have seen growing popularity for computing the properties of ordered systems with 1-3 elements [71], the corresponding pseudopotentials are not currently available, and the accuracy of these functionals for complex disordered systems with 5+ elements has yet to be assessed. Since the present work is intended as a starting point for scaling up to larger refractory alloy systems with increased chemical complexity, the PBE functional is utilized throughout this work, to ensure a uniform level of theory across all alloys.
@@
Initial benchmark studies were performed on bulk Mo, Nb, Ta, and W in order to determine appropriate computational parameters for the binary alloy calculations. We computed the lattice constants, bulk moduli, cohesive energies, elastic constants, and (100), (110), (111) surface energies for elemental Mo, W, Nb, and Ta, and compared our results with available experimental data, using three different pseudopotentials: the FHI norm-conserving pseudopotentials ($d$ valence electrons only) tabulated at the Abinit website [73], pseudo-dojo [68] NCPPs ($s$, $p$ and $d$ valence electrons) available at [74], and projector augmented-wave (PAW) pseudopotentials [75] ($s$, $p$ and $d$ valence electrons), tabulated at the Abinit website [73]. Computed properties using the FHI NCPPs showed an average deviation of $\approx 1\%$ compared to available experiments, across all computed properties for Mo, Nb and Ta, but an average deviation of $\approx 9\%$ for W. For the pseudo-dojo NCPPs and PAW-based pseudopotentials, computed properties agreed well with experiment, with an average deviation of $\approx 1.5\%$ for all elements and across all properties. Detailed results are reported in [19]. Given the comparable performance of the pseudo-dojo NCPP and PAW pseudopotentials, and since DFPT is not yet implemented for PAW-based pseudopotentials in Abinit, the pseudo-dojo NCPPs were used throughout this work.
@@ 3 Results
Table 1: Composition-dependent Voigt-averaged rigid-ion and relaxed-ion elastic constants for the binary alloys (Eq. (2.3), with standard deviations ($\pm$).
@@ 3.4 Elastic anisotropy as a function of composition
To further analyze the extent of elastic tensor property anisotropy in the refractory binary alloys, 3D plots of the orientation-dependent $G$, $E$ and $\nu$, for three compositions each, were generated using the ELATE tool [103]. The results for MoNb and WTa are shown in Figs. 4 and 5; the results for MoTa, MoW, NbTa, and WNb can be found in the Appendix. The axes of the 3D plots in the figures correspond to the crystal orientations [1 0 0]delimited-[]100[1\,0\,0][ 1 0 0 ], [0 1 0]delimited-[]010[0\,1\,0][ 0 1 0 ], [0 0 1]delimited-[]001[0\,0\,1][ 0 0 1 ] in the $x$, $y$, and $z$ directions, respectively. The maximum and minimum values of $G$ and $\nu$ are represented as transparent blue and solid green surfaces. The extent of anisotropy is seen to be greatest for MoNb. For the Young’s modulus $E$, 1 0 0delimited-⟨⟩100$\langle$ 1\,0\,0$\rangle$ 1 0 0 ⟩ is the stiffest direction and 1 1 1delimited-⟨⟩111$\langle$ $\,1\rangle$ 1 1 ⟩ is the most compliant direction in all compositions of the binary alloys except for $\rm{Nb_{0.25}Ta_{0.75}}$ and $\rm{W_{0.25}Ta_{0.75}}$, which have nearly perfect spherical shapes, corresponding to elastic isotropy. $\rm{Mo_{0.25}Nb_{0.75}}$, $\rm{W_{0.25}Nb_{0.75}}$ and $\rm{Nb_{0.75}Ta_{0.25}}$ deviate strongly away from spherical, indicating high elastic anisotropy. Interestingly, WTa exhibits near-isotropic behavior at all compositions as seen in Fig. 5. Similar trends are observed for $G$ and $\nu$. The patterns observed in the 3D plots are consistent with the reduced metrics $A_{E}$ and $A_{G}$ in Table 2. For $A_{E}$, all binaries showed near-isotropic behavior with deviations away from 1 of $\leq 1\%$ except for $\rm{Mo_{0.25}Nb_{0.75}}$, $\rm{Mo_{0.75}Nb_{0.25}}$ and $\rm{W_{0.25}Nb_{0.75}}$, where the deviations are $>3\%$. For $A_{G}$, all binaries exhibited near-isotropic behavior (deviations of $<1\%$), except for $\rm{Mo_{0.25}Nb_{0.75}}$, $\rm{Mo_{0.75}Nb_{0.25}}$, $\rm{Nb_{0.75}Ta_{0.25}}$ and $\rm{W_{0.25}Nb_{0.75}}$, which exhibit greater deviations away from unity. For $A_{G}^{R}$, all binaries showed near-isotropic behavior (deviations of $<1\%$), except for $\rm{Mo_{0.25}Nb_{0.75}}$, $\rm{Mo_{0.25}Ta_{0.75}}$, $\rm{Mo_{0.5}Ta_{0.5}}$, $\rm{Nb_{0.25}Ta_{0.75}}$ and $\rm{Nb_{0.75}Ta_{0.25}}$.

### v3

@@ Abstract
The elastic tensor provides valuable insight into the mechanical behavior of a material with lattice strain, such as disordered binary alloys. Traditional stress-strain methods have made it possible to compute elastic constants for high-throughput databases of ordered structures, as well as individually tailored alloy compositions. However, this approach depends on predetermined or iteratively-chosen strain tensors. This poses a significant challenge for systematic, composition-dependent studies of disordered materials requiring structural models with randomized atomic positions and possessing low symmetry. Density functional perturbation theory (DFPT) based on density functional theory provides a compelling alternative to stress-strain methods: it allows for an unbiased determination of the elastic tensor, as well as access to local field data derived from the underlying general response function framework. Despite its intrinsic flexibility and efficiency, DFPT has seen limited application to the study of disordered systems. At the same time, there is a growing need for expanded quantum mechanical data to improve predictive modeling of complex disordered material properties. Here we present results for the rigid-ion and relaxed-ion elastic tensors computed using DFPT, for a comprehensive set of structural refractory body-centered cubic (BCC) binary alloys of molybdenum (Mo), niobium (Nb), tantalum (Ta), and tungsten (W). For the first time, we have mapped the quantum-driven heterogeneity in elastic constants and shear modulus, and associated relaxation fields at each disordered structure lattice site, by computing the force response internal strain tensor ($\Lambda$) and displacement response internal strain tensors ($\Gamma$). Derived properties—the bulk modulus ($B$), shear modulus ($G$), Young’s modulus ($E$), Poisson’s ratio ($\nu$), Pugh’s ratio ($B/G$), Cauchy pressure and elastic anisotropy—are reported as a function of composition for all refractory binaries. The DFPT-computed mechanical properties data for the refractory binary alloys at systematically-varied Mo, Nb, Ta, and W compositions are in excellent agreement with available experimental data.
@@
refractory alloy, density functional theory, compositionally-complex alloys, elastic properties,
DFPT, property anisotropy
@@ 1 Introduction
The elastic tensor [1] and its elastic constant components give a complete description, in terms of deformation, of material response under external stress within the elastic limit. The elastic tensor can therefore serve as a valuable quantitative probe of elastic strain effects associated with changes in chemical composition, and their impact on mechanical [2, 3, 4, 5, 6], electronic [7, 8], thermodynamic [9, 10], and structural [11, 12, 13] properties in these materials.
@@
Refractory alloys show great potential to meet the demands for high-temperature materials due to their high melting points and ability to maintain mechanical properties at temperatures higher than current Ni-based superalloys [14]. The refractory elements Nb, Ta, Mo, and W of groups V and VI comprise the BCC structural refractory metals family, characterized by high mechanical strength and melting temperatures [15]. Their structural refractory binary alloys MoNb, MoTa, MoW, NbTa, WNb and WTa form solid solutions of BCC crystal structure across their entire compositional space [16], and form an attractive starting point for designing more complex disordered alloys. However, elastic constant measurements are typically reported only for a limited number of alloy classes and compositions, and experiments often exhibit substantial disagreement even between values measured using the same technique and on the same material [17, 18]. Multiple factors can influence differences in reported experimental values, such as impurities and measurement temperature [19, 20, 21, 22]. Previous efforts have focused on the computational determination of elastic constants using density functional theory (DFT) for selected refractory binary alloys, using stress-strain methods [23, 24, 25, 26, 27]. While these investigations have provided valuable insights, they were limited to a few binary systems or computed for a narrow compositional range. The computational methodologies employed in these studies were primarily validated using experimental data for pure elements. Since the electronic interactions in alloy systems are inherently more complex than in their elemental counterparts, the direct application of methods validated for pure elements to alloy systems may introduce significant uncertainties, potentially compromising the accuracy and practical relevance of the predicted elastic properties for these alloys. These challenges are compounded when modeling disordered materials [28]. Here, we establish and validate a first-principles modeling framework based on DFT [29, 30] for performing computationally-efficient studies of elastic tensor and derived properties of disordered alloys, validated for systematically-varied Mo, Nb, Ta, and W binary alloy compositions. DFT, as implemented in state-of-the-art plane wave pseudopotential periodic electronic structure codes such as Abinit [31, 32], provides direct access to the computation of energies and stresses for obtaining the elastic tensors of the materials, with only atomic structure and nuclear charges as inputs.
@@
The principal quantity to be considered is the elastic tensor and its elastic constant components. The stress-strain relation is given by:
@@
where $\bm{\sigma}$ is the $6\times 1$ stress tensor, the $C_{ij}$ are the elastic constants, and $\bm{s}$ is the $6\times 1$ strain tensor in Voigt notation [33].
In the traditional stress-strain approach, the elastic constants $C_{ij}$ are obtained by computing either the total energy or the stress of a crystal at a series of preselected strains [19, 34, 35, 36, 37]. The selection of strain sets can significantly influence the computed elastic constants, with variations exceeding 15% [38] in some cases. To mitigate this problem, high-throughput databases of ordered materials, such as the Materials Project [39], validate computed elastic constants by comparing results obtained from multiple strain sets [38]. In disordered alloys, such as the refractory binary alloys that are the focus of the present work, this approach is not practical due to the lack of crystal symmetry caused by intrinsic lattice strain.
An alternative and more general approach that avoids the need for tailored strains is based on analytical differentiation of the energy using linear response theory [40, 41, 42, 43]. In this method, the elastic tensor is obtained by expanding the total energy with respect to infinitesimal perturbation of the system using density functional perturbation theory (DFPT) [44].
DFPT has a number of important advantages in the present context: (i) a preselected list of strains as a function of material and composition is not required, which avoids the potential generation of unphysical structures for unexplored alloys; (ii) it is generally applicable to all types of complex alloy systems, including those with large unit cells or reduced crystallographic symmetry, making it particularly well-suited for studying disordered or low-symmetry alloys; (iii) in contrast to the stress-strain method, determining the DFPT elastic tensor requires only a single calculation rather than a series of strain applications; (iv) because DFPT is couched in terms of a response function framework, it provides access to local field information that cannot be obtained from the manual application of a stress-strain approach. As shown below, such maps provide quantitative measures of the microscopic elastic response of the binary alloys, for potential input to data-driven predictive models of materials properties.
DFPT for the elastic tensor was initially implemented for norm-conserving pseudopotentials (NCPPs) [45] and later extended to the projector augmented-wave (PAW) method by Martin et al. [46]. DFPT has been successfully demonstrated using both NCPPs [47, 48, 49] and PAW pseudopotentials [50, 51, 52, 53, 54, 55]. Despite its intrinsic advantages, DFPT for elastic properties has been applied to only a limited number of structural alloys [56]. This work represents the first demonstration of DFPT for computing the elastic properties of the structural refractory disordered binary alloys of Mo, Nb, Ta and W.
@@
Within the DFPT framework, the elastic tensor can be evaluated as the rigid-ion elastic tensor, $C_{ij}^{rigid}$, or as the relaxed-ion elastic tensor $C_{ij}^{relaxed}$ [40, 41, 42, 43]. $C_{ij}^{rigid}$ is computed with the relaxed initial structure, and strain as the perturbation; $C_{ij}^{relaxed}$ is a correction to the rigid-ion elastic tensor, with both strain and atomic displacements (nuclear relaxations) as the perturbations. Both $C_{ij}^{rigid}$ and $C_{ij}^{relaxed}$ provide essential information about the elastic response of a material. The relaxed-ion elastic tensor is the quantity that will correspond most closely to experiment.
@@
Comparison of the rigid-ion and relaxed-ion results allows an assessment of the accuracy of the rigid-ion approximation, as well as enabling the mapping of the relaxation field of a disordered alloy. We define the nuclear relaxation elastic tensor $C_{ij}^{\text{nuclear}}$ as the difference between the corresponding rigid-ion and relaxed-ion quantities [43]:
@@
In the present work, we utilize DFT+DFPT to compute the single-crystal, relaxed-structure elastic tensors of the six BCC refractory binary alloys of (Mo, Nb, Ta, W)—MoNb, MoTa, MoW, NbTa, MoTa, and MoW—as a function of composition. The disordered alloys are modeled as supercells constructed using the special quasi-random structures (SQS) method [58]. The relaxed disordered structures are then used to compute the rigid-ion and relaxed-ion elastic tensors based on DFPT. We assess the effectiveness of DFPT in providing accurate elastic tensors through detailed comparison of elastic constants with available experimental data. The composition-dependent calculations serve as an important test of the ability of the SQS method to accurately model disorder in these alloys. Mechanical properties $B$, $G$, and $E$, the elastic anisotropy and $B/G$ and $C^{\prime}$ derived from the computed elastic moduli are reported as a function of composition for all refractory binaries. Computed relaxation fields for the Voigt-averaged elastic constants and shear modulus are presented for all binaries and discussed in terms of constituent element chemistry. The in-depth understanding of trends in elastic properties and anisotropy in the disordered refractory binary alloys gained in this work through the decomposition of elastic constants into electronic and nuclear contributions is expected to contribute to new microscopic alloy computational strategies for designing refractory alloys for high-temperature applications.
@@ 2 Computational Methods
We used the open-source, periodic plane wave pseudopotential electronic structure code Abinit [31, 32] for all DFT calculations, employing the Perdew, Burke, and Ernzerhof (PBE) generalized gradient approximation (GGA) exchange-correlation functional [59], with corresponding PBE norm-conserving pseudopotentials [60]. We note that while meta-GGA methods such as r2SCAN [61] and parameterized DFT+$U$ [62] approaches have seen growing popularity for computing the properties of ordered systems with 1–3 elements [63], the corresponding pseudopotentials are not currently available, and the accuracy of these functionals for complex disordered systems with 5+ elements has yet to be assessed. Since the present work is intended as a starting point for scaling up to larger refractory alloy systems with increased chemical complexity, the PBE functional is utilized throughout this work, to ensure a uniform level of theory across all alloys.
@@
Initial benchmark studies were performed on bulk Mo, Nb, Ta, and W in order to determine appropriate computational parameters for the binary alloy calculations. We computed the lattice constants, bulk moduli, cohesive energies, elastic constants, and (100), (110), (111) surface energies for elemental Mo, W, Nb, and Ta, and compared our results with available experimental data, using three different pseudopotentials: the FHI norm-conserving pseudopotentials ($d$ valence electrons only) tabulated at the Abinit website [65], pseudo-dojo [60] NCPPs ($s$, $p$ and $d$ valence electrons) available at [66], and projector augmented-wave (PAW) pseudopotentials [67] ($s$, $p$ and $d$ valence electrons), tabulated at the Abinit website [65]. Computed properties using the FHI NCPPs showed an average deviation of $\approx 1\%$ compared to available experiments, across all computed properties for Mo, Nb and Ta, but an average deviation of $\approx 9\%$ for W. For the pseudo-dojo NCPPs and PAW-based pseudopotentials, computed properties agreed well with experiment, with an average deviation of $\approx 1.5\%$ for all elements and across all properties. Detailed results are reported in [68]. Given the comparable performance of the pseudo-dojo NCPP and PAW pseudopotentials, the pseudo-dojo NCPPs were used throughout this work. Working within the pseudo-dojo framework [60] also provides a foundation for the custom construction of pseudopotentials based on emerging functionals, for computationally-consistent modeling of alloys of growing complexity.
@@ 3 Results
Table 1: Composition-dependent Voigt-averaged rigid-ion and relaxed-ion elastic constants for the binary alloys (Eq. (2.3), with standard deviations ($\pm$).)
@@ 3.4 Elastic anisotropy as a function of composition
To further analyze the extent of elastic tensor property anisotropy in the refractory binary alloys, 3D plots of the orientation-dependent $G$, $E$ and $\nu$, for three compositions each, were generated using the ELATE tool [96]. The results for MoNb and WTa are shown in Figs. 4 and 5; the results for MoTa, MoW, NbTa, and WNb can be found in the Appendix. The axes of the 3D plots in the figures correspond to the crystal orientations [1 0 0]delimited-[]100[1\,0\,0][ 1 0 0 ], [0 1 0]delimited-[]010[0\,1\,0][ 0 1 0 ], [0 0 1]delimited-[]001[0\,0\,1][ 0 0 1 ] in the $x$, $y$, and $z$ directions, respectively. The maximum and minimum values of $G$ and $\nu$ are represented as transparent blue and solid green surfaces. The extent of anisotropy is seen to be greatest for MoNb. For the Young’s modulus $E$, [1 0 0]delimited-[]100[1\,0\,0][ 1 0 0 ] is the stiffest direction and [1 1 1]delimited-[]111[1\,1\,1][ 1 1 1 ] is the most compliant direction in all compositions of the binary alloys except for $\rm{Nb_{0.25}Ta_{0.75}}$ and $\rm{W_{0.25}Ta_{0.75}}$, which have nearly perfect spherical shapes, corresponding to elastic isotropy. $\rm{Mo_{0.25}Nb_{0.75}}$, $\rm{W_{0.25}Nb_{0.75}}$ and $\rm{Nb_{0.75}Ta_{0.25}}$ deviate strongly away from spherical, indicating high elastic anisotropy. Interestingly, WTa exhibits near-isotropic behavior at all compositions as seen in Fig. 5. Similar trends are observed for $G$ and $\nu$. The patterns observed in the 3D plots are consistent with the reduced metrics $A_{E}$ and $A_{G}$ in Table 2. For $A_{E}$, all binaries showed near-isotropic behavior with deviations away from 1 of $\leq 1\%$ except for $\rm{Mo_{0.25}Nb_{0.75}}$, $\rm{Mo_{0.75}Nb_{0.25}}$ and $\rm{W_{0.25}Nb_{0.75}}$, where the deviations are $>3\%$. For $A_{G}$, all binaries exhibited near-isotropic behavior (deviations of $<1\%$), except for $\rm{Mo_{0.25}Nb_{0.75}}$, $\rm{Mo_{0.75}Nb_{0.25}}$, $\rm{Nb_{0.75}Ta_{0.25}}$ and $\rm{W_{0.25}Nb_{0.75}}$, which exhibit greater deviations away from unity. For $A_{G}^{R}$, all binaries showed near-isotropic behavior (deviations of $<1\%$), except for $\rm{Mo_{0.25}Nb_{0.75}}$, $\rm{Mo_{0.25}Ta_{0.75}}$, $\rm{Mo_{0.5}Ta_{0.5}}$, $\rm{Nb_{0.25}Ta_{0.75}}$ and $\rm{Nb_{0.75}Ta_{0.25}}$.